Five apples and a pear cost $4.67. If an apple cost $0.45, how much does a pear cost?
Answer by Edwin McCravy(20054)


Using only basic math:

 $0.45    <-- cost of 1 apple
   x 5    <-- number of apples
 -----
 $2.25    <-- cost of 5 apples

 $4.67    <-- cost of 5 apples and 1 pear
-$2.25    <-- cost of 5 apples
------
 $2.42    <-- cost of 1 pear


Using algebra:


Five apples and a pear cost $4.67. 
  |    |     |     |    |     | 
  5    A     +     P    =    4.67

                 5A + P = 4.67

If an apple cost $0.45, 
        |    |     |
        A    =    0.45 

                      A = 0.45

how much does a pear cost?

Plug (0.45) for A into

                 5A + P = 4.67

            5(0.45) + P = 4.67

Multiply the 5 by the 0.45 and get 2.25 and
substitute that for 5(0.45)

               2.25 + P = 4.67

Subtract 2.25 from both sides:


               2.25 + P = 4.67
              -2.25      -2.25
           ------------------------
                      P = 2.42

Answer: $2.42

Edwin